5: P is a point on the parabola whose ordinate

equals its abscissa. A normal is drawn to the parabola at

P to meet it again at Q. If S is the focus of the parabola

then the product of the slopes of SP and SQ is-


1
Expert's answer
2021-11-07T16:48:21-0500

The answer to your question is provided in the image:


"\text{Since the abscissa and ordinate are equal for point P}\\
2at_1=at_1^2\\
t_1=2\\
\text{Slope PQ}\\
\dfrac{2at_1-2at_2}{at_1^2-at_2^2}=\dfrac{2}{t_1+t_2}\\
\dfrac{dy}{dx}=\dfrac{\frac{dy}{dt}}{\frac{dx}{dt}}=\dfrac{1}{t_1}=\dfrac{1}{2}\\
\text{Product of slope PQ and tangent is -1}\\
\dfrac{2}{t_1+t_2}\times{\dfrac{1}{2}}=-1\\
t_1+t_2=-1\\
\therefore\\
t_2=-3\\
\text{As we know slope of line between two points is deter mined by 
}\\
{y_1-y_2\over x_1-x_2}\\
\text{Slope SP}\\
\dfrac{0-2at_1}{a-at_1^2}=\dfrac{4}{3}\\
\text{Slope SQ}\\
\dfrac{0-2at_2}{a-at_2^2}=\dfrac{-3}{4}\\
\therefore\\
\text{The product of slope SQ and slope SP is -1}"